OpenStudy (anonymous):

PLease check my answer!! I will fan & medal :) i would say it's 1/3 or 1/6 Joey is having a party. He is deciding between two music themes, the 50's or the 80's. Also, he is trying to decide if he wants burgers, pizza, or hot dogs. He writes his options on slips of paper and places the music themes in one hat and the food choices in another. If he randomly pulls one option out of each hat, what is the probability that he pulls 50's and hot dogs?

OpenStudy (anonymous):

1/2 * 1/3 = 1/6 50's or 80's.....picking 50's, chances are 1/2 burgers,pizza,hot dogs....picking hotdogs, chances are 1/3

Miracrown (miracrown):

yeah 1/6... the reasoning is that we have a 1/2 shot at getting 50's music (since there are only two options) and we have a 1/3 shot at getting hot dogs (out of the 3 choices). 1/2*1/3 =1/6